OpenStudy (anonymous):

i will medal and fan! Express the ratio in simplest form. 24 : 18 A. 12 : 9 B. 4 : 3 C. 8 : 6 D. 6 : 1

OpenStudy (anonymous):

\[24 : 18 = \frac{24}{18} = \frac{6 \times 4}{6 \times 3}= \frac{4}{3}\]=4:3 Hence option B is the correct answer. @asfdgl